what is the difference between 126 1/4 and 78 2/3?
Answer by checkley77(12569)

126 1/4 and 78 2/3?
(126*4+1)/4=(504+1)/4=505/4=126.25
(78*3+2)/3=(234+2)/3=236/3=78.67
126.25-78.67=47.58 ans. for the difference.